A race car starting from rest accelerates uniformly at a rate of 4.90 meters per second2. What is the car’s speed after it has traveled 200. meters?
1
Expert's answer
2019-10-10T09:42:19-0400

We can use the formula:


"v_f^2-v_i^2=2as"

"v_f^2-0^2=2(4.9)(200)"

"v-f=44.3\frac{m}{s}"